# Non-stationary Diffusion For Probabilistic Time Series Forecasting

### v1

@@ 1 Introduction
Time series forecasting plays a key role in various fields such as traffic prediction (Ermagun & Levinson, 2018) and supply chain management (Chopra & Meindl, 2021). Given a historical multivariate series $\mathbf{X}$, general forecasting methods involve training an $f(\mathbf{X})$ to predict a future series $\mathbf{Y}$, which can be viewed as modeling the $\mathbb{E}[\mathbf{Y}|\mathbf{X}]$. Although recent research has demonstrated promising capabilities to model conditional expectations (Zhou et al., 2021; Wu et al., 2020; Zeng et al., 2023), effective decision-making, particularly in high-stakes fields like healthcare (Bertozzi et al., 2020) and finance (Li & Bastos, 2020), often requires accurately estimating the uncertainty underlying the data (Kendall & Gal, 2017). To address this problem, many recent studies have focused on probabilistic time series forecasting (Rasul et al., 2021; Li et al., 2024b), where the goal to estimate a distribution of possible future outcomes along with their associated uncertainties.
@@
Figure 1 illustrates an example from the ILI (influenza-like illness) dataset, with endpoint distributions (Left) and estimated uncertainty (Right) on different models: TimeGrad (Rasul et al., 2021), TMDM (Li et al., 2024b), and NsDiff (ours). In the upper part of Figure 1, TimeGrad (Rasul et al., 2021) employs the endpoint $\mathcal{N}(\mathbf{0},\mathbf{I})$, which fails to capture non-stationary characteristics. TMDM (Li et al., 2024a) uses $\mathcal{N}(f(\mathbf{X}),\mathbf{I})$ as endpoint, representing changing averages. On the test dataset (shown to the right of the red dashed line), where both the number of patients and the corresponding deviation increase, the performance differences are evident. TimeGrad fails to model both the underlying trends and deviations. In contrast, TMDM effectively captures the trends through its $f(\mathbf{X})$, but its stationary covariance $\mathbf{I}$ limits its ability to accurately estimate uncertainty, which is critical for the probabilistic time series forecasting.
To better address non-stationarity with changing uncertainty, we introduce Location-Scale Noise Model (LSNM) into DDPMs, which relaxes the traditional Additive Noise Model (ANM) by incorporating a contextually changing variance: $\mathbf{Y}=f(\mathbf{X})+\sqrt{g(\mathbf{X})}\boldsymbol{\epsilon}$, where $g(\mathbf{X})$ is an $\mathbf{X}$-dependent variance model. LSNM is capable of modeling both the contextual mean through $f(\mathbf{X})$ and the contextual uncertainty through $\sqrt{g(\mathbf{X})}$. In the special case where $g(\mathbf{X})\equiv 1$, this simplifies to the standard ANM. Building upon this more flexible and expressive assumption, we propose the Non-stationary Diffusion Model (NsDiff) framework, which provides an uncertainty-aware noise schedule for both forward and reverse diffusion processes. In summary, our contributions are as:
@@ 2.1 DDPM for Probabilistic Forecasting
Denoising Diffusion Probabilistic Models (DDPMs) have shown promising results in the probabilistic forecasting area (Tyralis & Papacharalampous, 2022). Rasul et al. (2021) introduce TimeGrad, an autoregressive diffusion model guided by a recurrent neural network hidden state. Tashiro et al. (2021) propose a masking strategy for training diffusion models, applicable to tasks like imputation and forecasting. Alcaraz & Strodthoff (2022) extend DDPMs with a structured space model to capture long-term dependencies. TimeDiff (Shen & Kwok, 2023) utilized future mixup and autoregressive initialization.  Li et al. (2022b) integrate multiscale denoising score matching to guide the diffusion process, ensuring generated series align with the target. DiffusionTS (Yuan & Qiao, 2024) trains the model to reconstruct the sample rather than noise, using a Fourier-based loss term. Kollovieh et al. (2024) propose a self-guiding strategy for time series generation and forecasting based on structured state-space models. These methods generally assume fixed endpoint variance, which is hard to model non-stationary time series.
@@ 2.2 Non-stationary Time Series Forecasting
To address non-stationarity, Li et al. (2022a) employ a domain-adaptation approach to predict data distributions, while Du et al. (2021) propose an adaptive RNN for distribution matching to mitigate non-stationary effects. Liu et al. (2022) introduce a non-stationary Transformer with de-stationary attention to account for non-stationary factors in self-attention. Wang et al. (2022) use global and local Koopman operators to capture patterns at different scales, and Liu et al. (2024a) apply Koopman operators to components identified via Fourier transforms. Other approaches decompose stationary and non-stationary parts, such as Ogasawara et al. (2010) with local normalization, and Passalis et al. (2019) with a learnable, instance-wise normalization. RevIN (Kim et al., 2021) addresses the distribution shift using reversible normalization, and recent works (Fan et al., 2023; Liu et al., 2024b) explore finer-grained trend modeling. Fourier transforms, closely linked with non-stationarity, are also been applied to tackle these issues (Fan et al., 2024; Ye et al., 2024). Despite these advances in time series forecasting, the non-stationary uncertainty in probabilistic forecasting remains largely unexplored.
@@ 4 Methodology
Figure 2: The outline of NsDiff. It integrates a LSNM-based endpoint and an uncertainty-aware noise schedule. During the training phase, a noise and variance estimator, $\xi_{\theta}$, is optimized to approximate the reverse process distribution. During inference, it samples from LSNM endpoint and use the estimated reverse distribution to iteratively denoise and generate the final prediction.
@@ 4.3 Pretraining $f_{\phi}$ and $g_{\psi}$
To train $f_{\psi}$, we follow prior work (Li et al., 2024a) and utilize the Non-stationary Transformer (Liu et al., 2022) as the backbone model. The training process is identical to that of standard supervised time series models (Zhou et al., 2021). For the training of $g_{\psi}(\mathbf{X})$, while the true distribution of point-wise time series data is intractable – since one can only sample once at a specific time point – we use a sliding window approach to extract the estimated ground truth variance, similar to reference (Kim et al., 2021; Liu et al., 2024b; Ye et al., 2024) Specifically, given time series label $\mathbf{Y}_{0}$ the estimated ground truth variance is defined as:
@@
thus, the training of $g_{\psi}(\mathbf{X})$ is formulated as a supervised task. In our implementation, we utilize a sliding stride of 1 and a window size of 96. The function $g_{\psi}$ is implemented as a three-layer MLP, with outputs passed through the softplus activation (Zheng et al., 2015) to ensure positivity. Further implementation details can be found in Appendix C.2.
@@ 4.5 Inference
Therefore, the solvability of the equation is governed by the noise level parameter $\beta_{t}$. Under the typical DDPM parameterization (Ho et al., 2020), where $\beta_{t}$ ranges from 0.0001 to 0.02, the coefficient on the right-hand side of Eq. 17 becomes sufficiently large, thereby ensuring the equation’s solvability. Experimentally, the approach exhibits consistent solvability across all datasets.
Hence, by solving the quadratic equation in Eq. 15, we can estimate the value of $\sigma_{\mathbf{Y}_{0}}$ during inference stage, the specific formula is given by:
@@
we provide more discussions in Appendix A.4.
@@ 5.1 Experiment Setup
Metrics: Following prior work (Li et al., 2024a), we use two probabilistic forecasting metrics: Quantile Interval Coverage Error (QICE) (Han et al., 2022) and Continuous Ranked Probability Score (CRPS) (Matheson & Winkler, 1976). For both metrics, smaller values indicate better performance. Detailed formula is provided in Appendix C.3. We provide point forecast results at Appendix B.
@@ A.5.2 Unit Variance
Appendix B Point Forecasting Results

### v2

@@ 1 Introduction
Time series forecasting plays a key role in various fields such as traffic prediction (Ermagun & Levinson, 2018) and supply chain management (Chopra & Meindl, 2021). Given a historical multivariate series $\mathbf{X}$, general forecasting methods involve training an $f(\mathbf{X})$ to predict a future series $\mathbf{Y}$, which can be viewed as modeling the $\mathbb{E}[\mathbf{Y}|\mathbf{X}]$. Although recent research has demonstrated promising capabilities to model conditional expectations (Zhou et al., 2021; Wu et al., 2020), effective decision-making, particularly in high-stakes fields like healthcare (Bertozzi et al., 2020) and finance (Li & Bastos, 2020), often requires accurately estimating the uncertainty underlying the data (Kendall & Gal, 2017; Xu et al., 2024). To address this problem, many recent studies have focused on probabilistic time series forecasting (Rasul et al., 2021; Chen et al., 2024; Li et al., 2024b), where the goal to estimate a distribution of possible future outcomes along with their associated uncertainties.
@@
Figure 1 illustrates an example from the ILI (influenza-like illness) dataset, with different endpoint distributions (Left) and estimated uncertainty (Right) on different models: TimeGrad (Rasul et al., 2021), TMDM (Li et al., 2024b), and NsDiff (ours). In the upper part of Figure 1, TimeGrad (Rasul et al., 2021) employs the endpoint $\mathcal{N}(\mathbf{0},\mathbf{I})$, which fails to capture non-stationary characteristics. TMDM (Li et al., 2024a) uses $\mathcal{N}(f(\mathbf{X}),\mathbf{I})$ as endpoint, representing changing averages. On the test dataset (shown to the right of the red dashed line), where both the number of patients and the corresponding deviation increase, the performance differences are evident. TimeGrad fails to model both the underlying trends and deviations. In contrast, TMDM effectively captures the trends through its $f(\mathbf{X})$, but its stationary covariance $\mathbf{I}$ limits its ability to accurately estimate uncertainty, which is critical for the probabilistic time series forecasting.
To better address non-stationarity with changing uncertainty, we introduce Location-Scale Noise Model (LSNM) into DDPMs, which relaxes the traditional Additive Noise Model (ANM) by incorporating a contextually changing variance: $\mathbf{Y}=f(\mathbf{X})+\sqrt{g(\mathbf{X})}\boldsymbol{\epsilon}$, where $g(\mathbf{X})$ is an $\mathbf{X}$-dependent variance model and $\boldsymbol{\epsilon}$ is a standard gaussian noise. LSNM is capable of modeling both the contextual mean through $f(\mathbf{X})$ and the contextual uncertainty through $\sqrt{g(\mathbf{X})}$. In the special case where $g(\mathbf{X})\equiv 1$, this simplifies to the standard ANM. Building upon this more flexible and expressive assumption, we propose the Non-stationary Diffusion Model (NsDiff) framework, which provides an uncertainty-aware noise schedule for both forward and reverse diffusion processes. In summary, our contributions are as:
@@ 2.1 DDPM for Probabilistic Forecasting
Denoising Diffusion Probabilistic Models (DDPMs) have shown promising results in the probabilistic forecasting area (Tyralis & Papacharalampous, 2022). Rasul et al. (2021) introduce TimeGrad, an autoregressive diffusion model guided by a recurrent neural network hidden state. Tashiro et al. (2021) propose a masking strategy for training diffusion models, applicable to tasks like imputation and forecasting. Alcaraz & Strodthoff (2022) extend DDPMs with a structured space model to capture long-term dependencies. TimeDiff (Shen & Kwok, 2023) utilized future mixup and autoregressive initialization.  Li et al. (2022b) integrate multiscale denoising score matching to guide the diffusion process, ensuring generated series align with the target. DiffusionTS (Yuan & Qiao, 2024) trains the model to reconstruct the sample rather than noise, using a Fourier-based loss term. Kollovieh et al. (2024) propose a self-guiding strategy for time series generation and forecasting based on structured state-space models. By leveraging the bridge-based model introduced by Shi et al. (2023), Chen et al. (2023) present a convergence analysis of the Schrödinger bridge algorithm and propose improvements to the diffusion process. However, these methods generally assume fixed endpoint variance, which is hard to model non-stationary time series.
@@ 2.2 Non-stationary Time Series Forecasting
To address non-stationarity, Li et al. (2022a) employ a domain-adaptation approach to predict data distributions, while Du et al. (2021) propose an adaptive RNN for distribution matching to mitigate non-stationary effects. Liu et al. (2022) introduce a non-stationary Transformer with de-stationary attention to account for non-stationary factors in self-attention. Wang et al. (2022) use global and local Koopman operators to capture patterns at different scales, and Liu et al. (2024a) apply Koopman operators to components identified via Fourier transforms. Other approaches decompose stationary and non-stationary parts, such as Ogasawara et al. (2010) with local normalization, and Passalis et al. (2019) with a learnable, instance-wise normalization. RevIN (Kim et al., 2021) addresses the distribution shift using reversible normalization, and recent works (Fan et al., 2023; Liu et al., 2024b) explore finer-grained trend modeling. Jiang et al. (2023) addresses non-stationarity in chaotic systems by preserving invariant measures to stabilize dynamical systems over time without relying on domain-specific priors. Fourier transforms, closely linked with non-stationarity, are also been applied to tackle these issues (Fan et al., 2024; Ye et al., 2024). Despite these advances in time series forecasting, the non-stationary uncertainty in probabilistic forecasting remains largely unexplored.
@@ 4 Methodology
Figure 2: The outline of NsDiff. It integrates a LSNM-based endpoint and an uncertainty-aware noise schedule. During the training phase, a noise and variance estimator, $\xi_{\theta}$, is optimized to approximate the reverse process distribution. To generate a sample, during inference, it samples from LSNM endpoint and use the estimated reverse distribution to iteratively denoise and generate the final prediction.
@@ 4.3 Pretraining $f_{\phi}$ and $g_{\psi}$
To train $f_{\psi}$, we follow prior work (Li et al., 2024a) and utilize the Non-stationary Transformer (Liu et al., 2022) as the backbone model. The training process is identical to that of standard supervised time series models (Zhou et al., 2021). For the training of $g_{\psi}(\mathbf{X})$, we use a sliding window approach to extract the estimated ground truth variance, similar to references (Kim et al., 2021; Liu et al., 2024b; Ye et al., 2024). Specifically, given time series label $\mathbf{Y}_{0}$ the estimated ground truth variance is defined as:
@@
thus, the training of $g_{\psi}(\mathbf{X})$ is formulated as a supervised task. In our implementation, we utilize a sliding stride of 1 and a window size of 96. The function $g_{\psi}$ is implemented as a three-layer MLP, with outputs passed through the softplus activation (Zheng et al., 2015) to ensure positivity. Further implementation details can be found in Appendix C.2 and we examine the necessity of pretraining in Appendix B.2.
@@ 4.5 Inference
Therefore, the solvability of the equation is governed by the noise level parameter $\beta_{t}$. Under the typical DDPM parameterization (Ho et al., 2020), where $\beta_{t}$ ranges from 0.0001 to 0.02, the coefficient on the right-hand side of Eq. 17 becomes sufficiently large, thereby ensuring the equation’s solvability. Hence, by solving the quadratic equation in Eq. 15, we can estimate the value of $\sigma_{\mathbf{Y}_{0}}$ during inference stage, the specific formula is given by Eq. 18:
@@
Experimentally, the approach exhibits consistent solvability across all datasets. We provide more discussions in Appendix A.4.
@@ 5.1 Experiment Setup
Metrics: Following prior work (Li et al., 2024a), we use two probabilistic forecasting metrics: Quantile Interval Coverage Error (QICE) (Han et al., 2022) and Continuous Ranked Probability Score (CRPS) (Matheson & Winkler, 1976). For both metrics, smaller values indicate better performance. Detailed formula is provided in Appendix C.3. We provide point forecast results at Appendix B.1.
@@ A.5.2 Unit Variance
Appendix B Other Results and Discussions
B.1 Point Forecast Results
@@
B.2 Effects of Pretraining
Although NsDiff follows the pretraining paradigm to stabilize training and achieve optimal performance, end-to-end training remains a viable alternative. Our experiments demonstrate that the networks can be trained jointly without significant performance degradation. For instance, when evaluating on the ETTh1 dataset with and without pretraining, we observe comparable results in the CRPS metric. We report the results in Table 8.
Table 8: The comparison between pretraining and end-to-end training, bold face indicate best result.
epoch pretrain end-to-end
1 0.4181 0.4407
2 0.4041 0.4227
3 0.3977 0.4045
4 0.3926 0.4004
5 0.3889 0.3868
6 0.3795 0.3873
As can be seen, although joint train experiences a slight performance degradation (1.86%), it still outperforms the previous state-of-the-art TMDM (0.452). However, compared to pretraining, co-training is slightly harder to converge. We will clarify this in the updated version.
B.3 Computation Efficiency
To analyse the computational efficiency of NsDiff, we compare the training and inference memory and time cost and performance, and report the results in Table 9.
Table 9: Computation efficiency comparison, bold face indicate best result.
Model Mem.Train(MB) Mem.Inference(MB) Tim.Train(ms) Tim.Inference(ms) CRPS QICE
TimeGrad 27.47 8.61 47.89 8319.29 0.606 6.731
CSDI 109.81 22.61 60.50 446.70 0.492 3.107
TimeDiff 15.66 3.40 33.93 238.78 0.465 14.931
DiffusionTS 65.03 79.23 94.51 8214.53 0.603 6.423
TMDM 221.58 213.46 33.26 237.37 0.452 2.821
NsDiff 68.20 57.75 32.13 208.07 0.392 1.470
As shown in Table 9, compared to previous SOTA TMDM, NsDiff achieves SOTA and has smaller memory costs and higher efficiency. This is because NsDiff does not introduce additional hidden variables and only adds a small number of basic operations. It should be noted that while NsDiff does not achieve the lowest memory cost, primarily due to its default use of the relatively heavy Non-stationary Transformer mean estimator (Liu et al., 2022), the NsDiff framework itself is model-agnostic. As such, the mean estimator can be replaced with a lighter conditional expectation forecasting model e.g. DLinear (Zeng et al., 2023) to optimize memory efficiency.